# Nagaoka Ferromagnetism in $3\times 3$ Arrays and Beyond

### v1

@@ IV.3 Removing the Center Site
When the loop sites are arranged evenly on a circle, each site position can be parameterized by an angle $\phi$ with $0\leq\phi\leq 2\pi$. Fixing one site on the loop at $\phi=0$, the correlation functions (CF) for the remaining electrons can be plotted against the angle. In Fig. 9, the CFs for a fixed spin up $CF(\uparrow)$ and spin down $CF(\downarrow)$ for three electrons on loops of various sizes are plotted against angle in the first and second columns. Here we display a GS when $s_{z}=1/2$. When system is in the IF regime for small $t/U$ ratio, the spin up/down CFs have peaks at angles $\phi=2\pi/3$ and $34\pi/$ 3, with the electrons as far away from each other as possible, and with a total spin of the ground state that is maximum at $S=3/2$. $CF(\uparrow)$ and $CF(\downarrow)$ are the same, showing that the two other electrons, regardless of spin, are correlated the same way to the fixed spin, as would be expected when interaction is dominant. After the transition when the $t/U$ ratio is large, the saturated ferromagnetism is lost, the two-peak structure disappears, the electrons behave qualitatively differently and the GS is no longer the maximum total spin. For a fixed spin-up electron, the other two electrons are spin up and spin down. With weak correlation, the other spin-up electron stays as far away as possible from the fixed spin-up electron. In contrast, when the spin-down electron is fixed, the remaining two electrons are spread almost uniformly around the entire ring with little correlation as expected when hopping is dominant.
@@
Figure 9: Dependence of the spin-up and spin-down correlation functions vs. angle for a loop with three electrons, $s_{z}=1/2$ and spin-up electron fixed at $\phi=0$. Different loop lengths are shown. The left column is for an IF GS below $t/U$ transition. The right column is for a GS above the transition where there is no IF.
@@
Note, the sign of $t$ matters for loops. For odd-numbered loops, the ferromagnetism for 3 electrons only occurs when $t<0$. There is no ferromagnetism for $t>0$ for any filling. For even numbered loops, the ferromagnetism for 3 electrons exists regardless of the sign of $t$. The even-numbered loops are bipartite and a transformation between $t$ and $-t$ can be done.
@@ V Summary and Discussion
In this work we have presented theoretical simulations done to probe Nagaoka ferromagnetism in arrays with different geometries. Starting from the $2\times 2$ plaquette, which was experimentally realized in [4], we have shown that perfect $3\times 3$ and $N\times N$ arrays up to $7\times 7$ exhibit NF for one-hole in a half-filled band, with a $t/U$ ratio for the onset of NF that decreases as the array size N increases. So far 2D dopant arrays have only been studied for the $3\times 3$ configuration. However, recent advances in fabrication of dopant structures gives significant promise for making highly engineered dopant arrays that will be an excellent testbed as an analog quantum simulator for this quantum magnetism.
@@
The existence of Nagaoka ferromagnetism or other forms of itinerant ferromagnetism crucially depends on the system geometry. One of the geometrical, necessary criteria is connectivity, not only in physical space, but also in the many-body spin-configuration space, which requires that any configuration of spins in the array can be transformed to any other configuration by a series of nearest neighbor hoppings without double occupancy. In $N\times N$ arrays, at least one hole is needed in a half-filled band for wavefunction connectivity at that filling. The $N\times N$ array has saturated ferromagnetism at this filling. A chain of $2\times 2$ square plaquettes with adjacent plaquettes connected at a common corner need at least two holes in a half-filled band to support wavefunction connectivity and supports saturated ferromagnetism for small $t/U$. In contrast, N site loops have wavefunction connectivity only for three-electron filling and saturated ferromagnetism for this filling.
A simple scaling argument shows that kinetic energy plays a critical role in determining the size dependence of the $t/U$ ratio at the onset of saturated ferromagnetism. In quasi two-dimensional systems, such as $N\times N$ arrays and chains of $2\times 2$ plaquettes, with ferromagnetism near half-filling, the critical $t/U$ decreases as the system size increases. A smaller hopping $t$ is needed to suppress ferromagnetism in larger systems because the kinetic energy increases more quickly that the repulsive energy as the size increases. In one-dimensional, N-site loops, with saturated ferromagnetism only for three electron filling, the critical $t/U$ increases as the loop size N increases and ferromagnetism can occur for larger $t$ and N increases. The scaling argument for small filling shows that the kinetic energy contribution becomes less important for larger loops, allowing the ferromagnetism to persist.
@@
The robustness of the saturated ferromagnetism to disorder in $N\times N$ arrays has been tested by considering $3\times 3$ arrays with one site removed. When a corner site is removed, saturated ferromagnetism is still possible for one hole in a half-filled band (which now has one less electron). When one of the edge sites is removed from the $3\times 3$ array saturated ferromagnetism is no longer possible because wavefunction connectivity is broken. When the center site is removed, the $3\times 3$ array becomes a loop. Saturated ferromagnetism is still possible but only for three electron filling, which is the only filling that has wavefunction connectivity. The transition between different realizations of ferromagnetism when a site is removed can be mapped out by varying the on-site energy of the site to be removed. The phase diagrams associated with these transitions will be discussed in a future publication.
@@
Interestingly, the first experiment to demonstrate Nagaoka ferromagnetism was for a $2\times 2$ plaquette. The $2\times 2$ plaquette is the smallest $N\times N$ array where one hole in a half filled band requires three electrons. However the $2\times 2$ array is also a loop that displays the three-electron ferromagnetism expected for loops. It is not clear whether ferromagnetism in $2\times 2$ plaquettes is best described as Nagaoka ferromagnetims or as three-electron loop ferromagnetism.
@@
Our results show that many forms of ferromagnetism are possible in complex, extended arrays. Both small-scale and extended dopant arrays should be a rich playground to explore quantum magnetism.
@@ Appendix A Scaling of the transition ratio with system size
There are two dramatic differences between Nagaoka ferromagnetism in $N\times N$ arrays and three-electron ferromagnetism in $N$-site loops. First, the band-filling necessary for saturated magnetism is related to the many-body wave function connectivity, which, in turn, is related to the system geometry. The geometries of the $N\times N$ array and the $N$-site are very different, leading to very different criteria for wave function connectivity. Nagaoka ferromagnetism in an $N\times N$ array occurs when there is one hole in a half-filled band. The number of electrons involved is $N\times N-1$. Three-electron ferromagnetism occurs in loops for three-electron filling, independent of loop size.
@@
The second essential difference is the scaling of the $t/U$ ratio at the ferromagnetic transition with the system size $N$. For $N\times N$ arrays, the $t/U$ ratio at the transition scales as $1/N^{3}$, as shown in Fig. 5b. As the array size $N$ increases, a larger $U$ is needed to transition to ferromagnetism, making it harder to reach ferromagnetism as $N$ increases. For $N$-site loops, Fig. 8 shows that the $t/U$ ratio at the transition scales as $N$. As the loop size increases, a smaller $U$ is needed for the transition to ferromagnetism, making it easier to for this transition to take place.
In this appendix we show how this dramatic difference in scaling can be understood based on simple arguments about how the difference in system hopping energy $\Delta T_{sys}(N)$ before and after the ferromagnetic transition and the corresponding difference in system electron-electron repulsion $\Delta U_{sys}(N)$ scale with the system size. To make the scaling argument, we assume that there is a critical ratio $R_{crit}^{sys}$ for $\Delta T_{sys}(N)/\Delta U_{sys}(N)$ that determines the relative strength of the hopping and interaction energies needed for the transition to ferromagnetism. We assume that $R_{crit}^{sys}$ may be different for loops and arrays, but does not depend on $N$. The usefulness of these two assumptions will be supported by the scaling relations that can be derived and by the success of the derived relations in explaining the opposite scaling with $N$ for loops and arrays.
A.1 Scaling of the number of participating electrons
For the $N\times N$ array with one hole in a half-filled band, the number of electrons contributing to the total hopping energy scales as $N^{2}$. The number of electron pairs contributing to the interaction energy scales as $N^{4}$. For the $N$-site loop the scaling is very different. For the three-electron ferromagnetism, three electron contribute to the hopping energy, independent of loop size. At most six electron-pairs contribute to the interaction energy, independent of loop size. This very different scaling for arrays and loops will explain most of the very different scaling for the $t/U$ ratio.
A.2 Scaling of the hopping energy difference
The transition to saturated ferromagnetism occurs once each ground state for total $s_{z}$ becomes the $s_{z}$-state with maximum total $S$. As shown in Figs. 2 and 5a, the ground state with maximum polarization transitions to the state with maximum $S$ at higher $t/U$ and the ground state with mimimum polarization transitions to the state with maximum $S$ last as $t/U$ is decrease. To estimate the scaling for the hopping energy difference we consider the transition of state with minimum polarization. This state will have nearly an equal number of up and down spins. For a nearly half-full band, the Fermi level for each spin will be at the middle of the band. States that can be rearranged in the ferromagnetic transition will be close to the middle of the band. The bandwidth of hopping energy in the array scales as $t$. There are $N^{2}$ states in the band for each spin. These states should have hopping energies that scale as $t/N^{2}$, especially near the band center. The number of electrons contributing to the transition should scale as $N^{2}$. The change in hopping energy of the array with one hole in a half-filled band should be independent of $N$.
For the loop, only three electrons contribute to the change in hopping energy and these states have hopping energies near the band edge, which scales as $t$. The total hopping energy for a filling with three electrons should scale as $t$.
A.3 Scaling of the interaction energy
The easiest way to establish the scaling is to consider the continuum limit for the array as a square hard-wall box. The interaction energy of a pair of electrons of opposite spin should be proportional to $U$ and the probability that the two spins overlap. In one dimension of the box, the probability that two plane-wave single-particle states overlap should scale as $1/L$ where $L$ is the length of the box side and $L$ scales as $N$. The probability of overlap for the square box will scale as $1/L^{2}$ when both dimensions are accounted for. The interaction per opposite-spin pair will scale as $U/L^{2}$ and the change in interaction energy $\Delta U_{array}(N)$ will scale as $UN^{2}$, where the number of opposite-spin pairs near the Fermi level scales as $N^{4}$.
For the loop, similar arguments show that the overlap along the loop scales as $1/L$. Unlike the square array, the loop is one dimensional so the scaling factor for the overlap differs for loops and arrays. The change in interaction energy per pair scales as $U/L$ and the total interaction energy for three electrons on the loop scales as $U/N$.
A.4 Scaling of the ratio $T_{sys}(N)/U_{sys}(N)$
Using the scaling relations for the change in hopping energy and the change in interaction energy for the array, the critical ratio for the array should scale as
$\Delta T_{array}(N)/\Delta U_{array}(N)\sim t/(UN^{2}).$
As a result, the $t/U$ ratio at the ferromagnetic transition for the array with one-hole in the half-filled band should scale as
$t/U\sim R_{crit}^{array}N^{2}.$
Using the scaling relations for the change in hopping energy and the change in interaction energy for the loop, the critical ratio for the loop should scale as
$\Delta T_{loop}(N)/\Delta U_{loop}(N)\sim(t)/(U/N).$
As a result, the $t/U$ ratio at the ferromagnetic transition for the loop with three electrons should scale as
$t/U\sim R_{crit}^{loop}/N.$
Most importantly, the scaling relations for $t/U$ show and explain the opposite behaviour for loops and arrays. For loops, $t/U$ scales inversely with N just as our numerical results show. In contrast, for arrays the scaling relation predicts quadratic scaling with $N$, slower scaling than the $N^{3}$ scaling in the numerical results. Three factors contribute to the different scaling. Most importantly, electrons near the middle of band contribute to the scaling for arrays. For loops, only a three electrons near the band edge contribute. For arrays there are extra factors of $N^{2}$ and $N^{4}$ that determine the scaling of hopping and interaction energy, For arrays, the hopping energy for electrons near the middle of the band has a $1/N^{2}$ scaling, while for loops, the hopping energy of three electrons near the band edge scales independently on $N$. Finally, the interaction energy per pair scales as $1/N^{4}$ in two-dimensional arrays, as compared to the $1/N^{2}$ scaling in one-dimensional loops.

### v2

@@ IV.3.4 Particle Correlation in Loop Ferromagnetism
Loop ferromagnetism can be further characterized by evaluating the two-particle correlations. When the loop sites are arranged evenly on a circle, each site position can be parameterized by an angle $\phi$ with $0\leq\phi\leq 2\pi$. Fixing one site on the loop at $\phi=0$, the correlation functions (CF) for the remaining electrons can be plotted against the angle. In Fig. 10, the CFs for a fixed spin up $CF(\uparrow)$ and spin down $CF(\downarrow)$ for three electrons on loops of various sizes are plotted against angle in the left column for a ferromagnetic ground state and in the right column for a nonferrromagnetic ground state. Here we display a GS when $s_{z}=1/2$. When the system is in the IF regime, shown here for a small $t/U$ ratio, the spin up/down CFs have peaks at angles $\phi=2\pi/3$ and $34\pi/$ 3, with the electrons as far away from each other as possible, and with a total spin of the ground state that is maximum at $S=3/2$. $CF(\uparrow)$ and $CF(\downarrow)$ are the same, showing that the two other electrons, regardless of spin, are correlated the same way to the fixed spin, as would be expected when interaction is dominant and the wavefunction is connected. When the $t/U$ ratio is large, the saturated ferromagnetism is lost, the two-peak structure disappears, the electrons behave qualitatively differently and the GS is no longer the maximum total spin. For a fixed spin-up electron, the other two electrons are spin up and spin down. With weak correlation, the other spin-up electron stays as far away as possible from the fixed spin-up electron, due to Pauli blockade, while the weakly correlated spin-down electron is more evenly spread around the loop. In contrast, when the spin-down electron is fixed, the remaining two electrons are spread almost uniformly around the entire loop with little correlation as expected when hopping is dominant.
@@
Figure 10: Dependence of the spin-up and spin-down correlation functions versus angle for a loop with three electrons, $s_{z}=1/2$ and the fixed electron-spin at $\phi=0$. Different loop lengths are shown. The left column is for a GS below the $t/U$ transition that is IF. The right column is for a GS above the transition where there is no IF.
IV.3.5 The Sign of t in Loop Ferromagnetism
@@
The sign of $t$ matters for loops. For odd-numbered loops, the ferromagnetism for three electrons only occurs when $t<0$ and for two electrons only for $t>0$. For even numbered loops, the ferromagnetism for three electrons exists regardless of the sign of $t$, and there is no ferromagnetism for any other filling. The even-numbered loops are bipartite and a transformation between $t$ and $-t$ can be done. This can also be explained with the exchange interaction coming from itinerant holes, when the exchange interaction is negative, the ground state is ferromagnetic. The specific combination of the sign of $t$, even/oddness of $N$ and electron filling determines the sign of the exchange interaction, and hence determines whether ferromagnetism exists.
@@ V Summary and Discussion
In this work we have presented theoretical simulations done to probe Nagaoka ferromagnetism in arrays with different geometries. Starting from the $2\times 2$ plaquette, which was experimentally realized in [4], we have shown that perfect $3\times 3$ and $N\times N$ arrays up to $8\times 8$ exhibit NF for one-hole in a half-filled band, with a $t/U$ ratio for the onset of NF that decreases as the array size N increases, with the transition scaling as $1/n^{3.5}$. So far, 2D dopant arrays have only been studied for the $3\times 3$ configuration. However, recent advances in fabrication of dopant structures give significant promise for making highly engineered dopant arrays that will be an excellent testbed as an analog quantum simulator for this quantum magnetism .[25] Extending coupled quantum dot systems to large, complex and extended arrays shows similar promise.
@@
The existence of Nagaoka ferromagnetism or other forms of itinerant ferromagnetism in small arrays crucially depends on the array geometry. One of the geometrical, necessary criteria is connectivity, not only in physical space, but also in the many-body spin-configuration space, which requires that any configuration of spins in the array can be transformed to any other configuration by a series of nearest neighbor hoppings without double occupancy. In $N\times N$ arrays, at least one hole is needed in a half-filled band for wavefunction connectivity at that filling. The $N\times N$ array has saturated ferromagnetism at this filling. A chain of $2\times 2$ square plaquettes with adjacent plaquettes connected at a common corner needs at least two holes in the half-filled band to support wavefunction connectivity and supports saturated ferromagnetism for small $t/U$. In contrast, $N$-site loops have wavefunction connectivity only for three-electron filling and saturated ferromagnetism for three-electron filling.
@@
The robustness of the saturated ferromagnetism to disorder in $N\times N$ arrays has been tested by considering $3\times 3$ arrays with one site removed. When a corner site is removed, saturated ferromagnetism is still possible for one hole in a half-filled band (which now has one less electron). When one of the edge sites is removed from the $3\times 3$ array, saturated ferromagnetism is no longer possible because wavefunction connectivity is broken. When the center site is removed, the $3\times 3$ array becomes a loop. Saturated ferromagnetism is possible but only for three-electron filling. Higher electron fillings don’t have wavefunction connectivity. The transition between different realizations of ferromagnetism when a site is removed can be mapped out by varying the on-site energy of the site to be removed. The phase diagrams associated with these transitions will be discussed in a future publication.
Electron filling plays an essential role in determining the type of ferromagnetism that can occur and the scaling of the transition $|t/U|$ that marks the crossover to ferromagnetism. Loop ferromagnetism occurs for the same, fixed, small electron filling for all loop sizes $N$. The ferromagnetic state occurs for three electron filling but not for more electrons. Nagaoka ferromagnetism in $N\times N$ occurs when there is one hole in a half-filled band. As one consequence the scaling of the transition $|t/U|$ is linear in $N$ for loops and $1/N^{3.5}$ in arrays. Loop ferromagnetism occurs for a broader range of $U$ as $N$ increases, while for NF, the ferromagnetic state occurs for a smaller range of $U$ as $N$ increases. Scaling arguments support and elucidate the scaling obtained from the computational studies. Three-electron loop ferromagnetism occurs when the kinetic energy costs of becoming ferromagnetic is compensated by the elimination of repulsive interaction in the ferromagnetic state. For NF, the transition out of the ferromagnetic state occurs when the hopping energy cost of the hole, partially localized by scattering from the minority spins is compensated by the exchange energy that is gained when hopping to create doubly occupied states is allowed at finite $U$.
@@
Interestingly, the first experiment to demonstrate Nagaoka ferromagnetism was for a $2\times 2$ plaquette. For $2\times 2$ plaquette, one hole in a half-filled band requires three electrons. However the $2\times 2$ array is also a loop that displays the three-electron ferromagnetism expected for loops. Ferromagnetism in $2\times 2$ plaquettes could be described as Nagaoka ferromagnetism or as three-electron loop ferromagnetism. These are very different, but it is not clear if one is a better characterization of the ferromagnetism of a $2\times 2$ plaquette.
@@
Our results show that many forms of ferromagnetism should be possible in arrays with various geometries. Both small-scale and extended dopant and quantum dot arrays should be a rich playground for quantum simulators to explore quantum magnetism.
@@ Appendix A Scaling of the transition ratio with system size
There are two dramatic differences between Nagaoka ferromagnetism in $N\times N$ arrays and three-electron ferromagnetism in $N$-site loops. First, the band-filling necessary for saturated magnetism is related to the many-body wave function connectivity, which, in turn, is related to the system geometry. The geometries of the $N\times N$ array and the $N$-site are very different, leading to very different criteria for wave function connectivity. In $N\times N$ arrays, at least one hole is needed in a half-filled band to have the wave function connectivity needed for saturated ferromagnetism. For one-hole doping, the electron number scales as $N^{2}$. Loop ferromagnetism occurs for three-electron filling, independent of loop size. Wave function connectivity in loops is only possible in loops for three-electron filling.
@@
The second essential difference is the scaling of the $t/U$ ratio at the ferromagnetic transition with the system size $N$. For $N\times N$ arrays, the $t/U$ ratio at the transition scales as $1/N^{3.5}$, as shown in Fig. 5b. As the array size $N$ increases, a larger $U$ is needed to transition to ferromagnetism, making it harder to reach ferromagnetism as $N$ increases. For $N$-site loops, Fig. 9 shows that the $t/U$ ratio at the transition scales as $N$. As the loop size increases, a smaller $U$ is needed for the transition to ferromagnetism, making it easier to for this transition to take place.
In this appendix we first show how the scaling for three-electron loop ferromagnetism can be understood based on simple arguments about how the difference in system hopping energy before and after the ferromagnetic transition and the corresponding difference in system electron-electron repulsion scale with the system size. Then we show how scaling for NF in an $N\times N$ array can be understood.
A.1 Scaling of the hopping energy in three-electron loops
Loop ferromagnetism (LF) occurs for three electron filling. The three electrons fill states with hopping energies near the band edge unless the loop is very short. Near the band edge, single-particle energies take the form $-a|t|+b|t|/N^{2}$, where $a$ defines the band edge, $b$ is the coefficient for the quadratic quantum confinement contribution and only the lowest order contribution in $1/N$ is kept. The total hopping energy for three electrons is $-3a|t|+b^{\prime}|t|/N^{2}$. The quantum confinement contribution for three electrons is determined by $b^{\prime}>0$. In LF the three electrons are fully polarized, have the same spin and must occupy higher energy states than in states with no loop ferromagnetism (nLF). As a consequence, $b^{\prime}_{LF}-b^{\prime}_{nLF}>0$ and the hopping energy increases when the three-electron ground state is the LF state.
A.2 Scaling of the interaction energy in three-electron loops
In the LF state, the three electrons have the same spin and there is no interaction energy. In contrast, in the nLF state, the state is only partially polarized and the interaction energy arises from the repulsive interaction $U$ between each pair of electrons with opposite spins that are on the same site. The probability that two electrons are on the same site in the loop scales as $1/N$. Thus the repulsive interaction energy scales as $n_{int}U/N$, where $n_{int}$ is the number of interacting pairs and is of order 1.
A.3 Scaling the transition from nLF to LF in three-electron loops
The crossover from nLF to LF occurs when the the additional hopping energy cost compensates the reduction of the interaction energy:
$(b^{\prime}_{LF}-b^{\prime}_{nLF})|t|/N^{2}=n_{int}U/N$
As a result of the scaling of the hopping and interaction energies, $|t|/U$ at the phase transition scales as $N$. This is the scaling obtained from diagonalization of the three-electron Hamiltonian for loops with different sizes $N$.
A.4 Using the states of the $N\times N$ array in the large $U$ limit to establish the scaling of the transition
To establish the scaling of the transition point in $N\times N$ arrays, it is easiest to start in the infinite $U$ limit and then determine how the energy levels shift as $U$ becomes finite and moves to the transition point. The dependence on $|U/t|$ of the energy-level structure for a $3\times 3$ array with one hole in a half-filled band is shown in Fig. 11 with the transition from no Nagaoka ferromagnetism (nNF) to Nagaoka ferromagnetism (NF) occurring in the region with the higher density of points. Energy levels are shown for the spin configuration with seven majority spins and one minority spin. The level structure for other spin configurations is similar, except with more levels. The lowest energy levels in the infinite $U$ limit are independent of $U$ because these are states with no doubly occupied sites. States with one doubly occupied site are highlighted in blue and can be identified as the lowest energy states with energies that increase linearly with $|U/t|$. Higher energy states with more than one doubly occupied site are not shown. The level crossing that determines the nNF to NF transition occurs in the levels for states with no doubly occupied sites.
Figure 11: Energy levels of a $3\times 3$ array with one hole in a half-filled band with seven majority spins and one minority spin. (a) Full spectrum, dependence on $|U/t|$. The region with a high-density of points is the region where the crossover between NF and nNF occurs. The blue-shaded region highlights the energies of states with one doubly-occupied site. (b) Blow up of the region near the transition.
The level structure is easiest to understand in the large $U$ limit with a simple structure above and near the crossover. The levels are easy to follow as $U$ becomes finite and moves toward the crossover regime. Below the crossover regime, the levels exhibit many crossings making it hard to follow the levels as $U$ varies. To understand the scaling of the transition $|U/t|$ on $N$, we will follow the levels from the infinite $U$ limit down to the crossover regime.
In the large $|U/t|$ limit, it is more instructive to work with the Hamiltonian $\mathcal{H}$ scaled by $U$:
$\mathcal{H}_{sc}=\mathcal{H}/U=(t/U)\sum_{<i,j>,\sigma}c^{{\dagger}}_{i,\sigma}c_{j,\sigma}+\sum_{i}n_{i,\uparrow}n_{i,\downarrow.}$ (3)
The first term is the hopping term and is a small contribution that can be treated as a perturbation. The second term, independent of $U$, is the large contribution that determines the form of the states in the large $|U/t|$ limit. In the infinite $U$ limit, all states with $n_{do}$ doubly occupied sites have an energy of $n_{do}$. The states with no doubly occupied sites form a degenerate ground state subspace. This degeneracy can be broken by using degenerate zero-order perturbation theory and diagonalizing the hopping term in the subspace of states with no double occupancy.
Nagaoka ferromagnetism occurs for one hole in a half-filled band. When all spins are polarized in the same direction, one-hole doping corresponds to one hole in the completely filled band of majority-spin electrons. In the following we will describe states by the number of holes in the filled majority-spin band plus the number of minority-spin electrons. For simplicity we describe such a state by the number of holes and electrons. To have the band filling needed for NF, we must have one more hole than than the number of electrons. In the subspace where there are no doubly occupied sites, each electron must sit on the same site with one of the holes. There will be a bound electron-hole pair for each electron and one extra free hole. An electron can’t move away from the hole that it is paired with unless it can hop directly to the site occupied by the free hole. The holes bound to electrons can’t move away from the electron they are paired with because that would create a doubly occupied site. Only the free hole can move by hopping.
A.5 Scaling of the hopping energy for $N\times N$ arrays in the infinite $U$ limit
Using the model just described, we now analyze the energy-level structure of states with band filling needed for NF in the infinite $U$ limit. The lowest energy state in the subspace with $n_{do}=0$, should be the fully polarized state with one hole and no electrons. In this state, the hole moves freely in the $N\times N$ array with the motion determined by the hopping term in $\mathcal{H}_{sc}$. The energy levels are those of a free particle hopping in an $N\times N$ array. The lowest level at the band edge has an energy that scales with $-|t|/U$ (as $-|t|$ for the Hamiltonian without scaling). The excited levels close to the band edge are shifted to higher energy as $|t|/UN^{2}$ by quantum confinement effects and levels further away are split by $|t|/UN$ due to the density of levels. As a consequence, the excited levels close to the band edge should have a splitting from the band edge that scales between $|t|/UN$ and $|t|/UN^{2}$ (between $|t|/N$ and $|t|/N^{2}$ for the Hamiltonian without scaling).
The simplest states with partial polarization would have two holes and one electron, with one of the holes bound to the electron on the same site. The bound pair can’t move, unless it exchanges with the free hole, which is only possible when the hole and bound pair are nearest neighbors. The unbound hole is nearly free, able to move through the lattice except where the bound pair is. The bound pair acts to partially localize the nearly free hole, increasing the energy of the nearly free hole. As a consequence the fully polarized ground state, with no bound pairs, is the lowest energy state in the infinite $U$ limit. There are on the order of $N^{2}$ unique positions where the bound pair can sit. The existing single-hole energy levels should be further split into different levels for each partially localized state with level splittings scaling as $|t|/N^{4}$ near the band edge and $|t|/N^{3}$ away from the band edge. As a consequence, the excited levels close to the band edge should have a splitting from the band edge that scales between $|t|/UN^{3}$ and $|t|/UN^{4}$ (between $|t|/N^{3}$ and $|t|/N^{4}$ for the Hamiltonian without scaling).
More complicated states with less polarization would have one nearly free hole and multiple bound electron-hole pairs. The same arguments about the bound pairs localizing the nearly-free hole should apply. As a consequence, the excited levels close to the band edge should have a splitting from the band edge that scales between $|t|/UN^{3}$ and $|t|/UN^{4}$ (between $|t|/N^{3}$ and $|t|/N^{4}$ for the Hamiltonian without scaling). In addition, the same arguments imply that the fully polarized ground state has lower energy than any partially polarized state in the infinite $U$ limit.
A.6 Scaling the exchange interaction for $N\times N$ arrays for finite repulsive interaction
These arguments define the level structure and the scaling of the level splitting between the fully polarized ground state and the excited states with no doubly occupied sites in the infinite $U$ limit. We now need to determine how this leveling splitting changes as $U$ becomes finite. To do this we need to determine the second-order change in energy of the excited states due to hopping term in $\mathcal{H}_{sc}$. The energy of the fully polarized ground state (for large $U$) is independent of $U$ because there in no same-site occupation and no contribution from $U$ in this case. For excited states the second-order change arises from hopping to a single doubly-occupied site. For $\mathcal{H}_{sc}$, this second-order energy shift scales as $-(t/U)^{2}$ (for $\mathcal{H}$, this second-order energy shift scales as $-t^{2}/U$. This second order change due to exchange coupling to the states with one doubly-occupied site is revealed in Fig.11 as level repulsion between states with one doubly-occupied site and the low-energy excited states with no doubly-occupied sites.
A.7 Scaling the transition from nNF to NF for $N\times N$ arrays
The transition from NF to nNF as $U$ decreases from the infinite $U$ limit occurs when the energy shift from the exchange coupling to the states with a doubly-occupied site closes the gap between the excited state and the fully polarized ground state. Using $\mathcal{H}_{sc}$, the second-order exchange $-(t/U)^{2}$ closes the gap which scales between $|t|/UN^{3}$ and $|t|/UN^{4}$. As a consequence, the scaling of $|t|/U$ at the transition should be between $1/N^{4}$ and $1/N^{3}$. This is consistent with our calculations for different $N$ which show a $1/N^{3.5}$ scaling.
Appendix B Other Ferromagnetism
B.1 NF in Non-squares Arrays
We find that NF exists in rectangular and hexagonal arrays, when the squares or hexagons are connected by sharing a side. See the following Table 3.
Geometry Sketch NF/IF #$\#e$ Transition Details
2x3 (2 side-sharing sq) [Uncaptioned image] NF 5 $0.019$ 1-hole NF
2x4 (3 side-sharing sq) [Uncaptioned image] NF 7 $0.012$ 1-hole NF
2x5 (4 side-sharing sq) [Uncaptioned image] NF 9 $0.008$ 1-hole NF
2 side-sharing hexagon [Uncaptioned image] NF IF 9 5 $0.0022$ $0.0405$ 1-hole NF 5e-IF
Table 3: Summary table for other geometries with NF, their sketch, number of electrons, and $t/U$ transition point.
For hexagonal arrays, the one hole ferromagnetism is NF, while the 5 electron itinerant ferromagnetism comes when the structure is two loops sharing a side. This 5-electron ferromagnetism can be extended to other numbers of sites on the loop, we are able to calculate up to two 18-site loops sharing a side.
B.2 Two-hole IF in plaquette sharing corners
The two-hole IF extends to larger chains of squares, see following Table 4.
Geometry Sketch NF/IF #$\#e$ Transition Details
no opposite corner (2 corner-sharing sq) [Uncaptioned image] IF 5 $0.0127$ 2-hole IF
3 corner-sharing sq [Uncaptioned image] IF 8 $0.0015$ 2-hole IF
4 corner-sharing sq [Uncaptioned image] IF 11 $0.0011$ 2-hole IF
5 corner-sharing sq [Uncaptioned image] IF 14 $0.0002$ 2-hole IF
Table 4: Summary table for two-hole IF, their sketch, number of electrons, and transition point if applicable.
These chains of plaquettes sharing corners need not be in a straight chain with the connecting corners along a common diagonal as shown in Table 4. For example, the connecting corners could be along a line made from a side of each plaquette. More complicated geometries for chains of squares also show IF if certain symmetry requirements are satisfied. This will be discussed in a future publication.
B.3 Robustness to extended Coulomb interaction
In actual experiments with dopant arrays [21], long-range Coulomb interactions ($V$) exist, and NF is robust for nearest neighbor Coulomb interaction $V_{nn}$ up to $V_{nn}/U\approx 0.017$ in $3\times 3$ arrays. This has been checked with calculations not reported here.